OpenStudy (anonymous):

A ball of mass 0.2 kg falls through air with an acceleration of 6 m/sec^2. Calculate the air drag on the ball. I have the problem about the word AIR DRAG, whats it actually?

OpenStudy (anonymous):

write down all the forces acting on the freely-falling body....ALL also draw adiagram if u wish

OpenStudy (anonymous):

then u will figure out wat air drag is

OpenStudy (mos1635):

AIR DRAG= force on body due to resistance of air. this force depence on shape of body, its velocity and density of air...

OpenStudy (anonymous):

net force acting on body=sum of all operating forces air drag is a force all given above is only u need to understand the problem

OpenStudy (anonymous):

ok 0.2 x 10 - F(air drag) = 0.2 x 6 => F= 2- 1.2 = 0.8 N

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

yes u are

OpenStudy (anonymous):

k tnx